I have another word problem that I need some help with. A vending machine contains nickels and dimes worth $14.50. There are 95 more nickels than dimes. How many of each coin is there? Also there is another word problem that has to do with an airplain flying against the wind. An airplane flew for 4hr. with a 20-km/hr. tail wind. The return flight against the same wind took 5hr. Find the speed of the plane in still air.
Answer by stanbon(75887)

A vending machine contains nickels and dimes worth $14.50. There are 95 more nickels than dimes. How many of each coin is there?
Let # of dimes be "x"; Value of these is 10x cents
Number of nickels is "x+95" ; Value of these is 5(x+95)=5x+475 cents
EQUATION:
value + value = value
10x+5x+475=1450 cents
15x=975
x=65 (number of dimes)
x+95=160 (number of nickels)
---------------------
Also there is another word problem that has to do with an airplain flying against the wind. An airplane flew for 4hr. with a 20-km/hr. tail wind. The return flight against the same wind took 5hr. Find the speed of the plane in
still air.
Down-wind DATA:
Whoops! Have to go
Cheers,
Stan H.